OpenStudy (anonymous):

fine the equation of the line containing the given pair of points (-5,-1)(-9,-4)

OpenStudy (dumbcow):

slope = (-1+4)/(-5+9) = 3/4 using point-slope form y +1 = 3/4(x+5) y = 3/4x + 11/4